OpenStudy (anonymous):

factor: 6x^2 + x = 5

Parth (parthkohli):

Try to get 0 in the RHS. Achieve that by subtracting 5 from both sides.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow 6x^2 + x - 5 = 5 - 5 }\]

Parth (parthkohli):

Now try to factor.

Parth (parthkohli):

Remember that -5 * 6 = -30.